A glider is undergoing simple harmonic motion on a frictionless air track. When it is at position x = 0.30 m, its kinetic energy is half of its potential energy, K=1/2U. What is the amplitude of its simple harmonic motion?
1
Expert's answer
2019-11-13T09:26:01-0500

The potential energy of the harmonic oscillator is "U = \frac{m \omega^2 x^2}{2}", and total energy is "E = T + U = \frac{m \omega^2 A^2}{2}", where "A" is the amplitude, "m" is the mass, "\omega" is the angular frequency. .

At "x = 0.3 m", the kinetic energy is "T = \frac{U}{2}", so at that moment "E = \frac{U}{2} + U = \frac{3 U}{2} = \frac{3}{2} \frac{m \omega^2 x^2}{2}", and according to the previous equation "E = \frac{m \omega^2 A^2}{2}". Hence, at "x = 0.3", "\frac{3}{2} \frac{m \omega^2 x^2}{2} = \frac{m \omega^2 A^2}{2}", from where "A^2 = \frac{3 x^2}{2}" and "A = \sqrt{\frac{3}{2}}x \approx 0.37 m".